# Absolute continuity of non-Gaussian and Gaussian multiplicative chaos measures

### v1

@@ 4. Generalizations and Open Questions
Here, we discuss some natural extensions and open questions raised by our work. To explain what our proof strategy can and can not handle, we frequently refer to the proof outline in Section 1.3. In particular, recall from there the “discretization step” and “coupling step.”
4.1. Absence of $\gamma\in(0,1]$
The immediate follow-up question to our paper is how to deal with $\gamma\in(0,1]$. As already detailed in Section 1.3, the only step of the proof which needs $\gamma>1$ is the coupling step using the Yurinskii coupling (Lemma 2.5). As mentioned in Remark 1.2, our result extends to $\gamma\in(\sqrt{2/3},\sqrt{2})$ under the additional assumption that $\mathbb{E}[(a_{1}^{(1)})^{3}]=0$: this can be seen from [new-yurinskii, Corollary 2.3], which is a strengthened, more general version of the Yurinskii coupling. From examining their proof, it seems plausible that assuming higher moments of $a_{1}^{(1)}$ match that of a Gaussian leads to a stronger coupling, which in particular would extend the range of $\gamma$ for which our result holds.
@@ 4.2. Critical chaos
The construction of critical multiplicative chaos has not yet appeared for the random Fourier series (1.1); indeed, it is left as an open question (for a more general class of non-Gaussian fields) in [Junnila, Section 3.3]. However, if it were constructed, we are hopeful that our methods would yield mutual absolute continuity with respect to a coupled critical GMC, in light of the fact that the number of thick points only decreases as $\gamma$ increases. Therefore, the coupling step becomes easier. The discretization step is not $\gamma$-dependent, so the difference comes from the proof of Proposition 2.7, which is achieved through the use of exponential moments alone.
@@ 4.3. Extension to general bounded domains
In this article, we considered the space $D=[0,1]$ for convenience. However, given that the range of $\gamma$ for which our proof breaks down excludes the $L^{2}$-regime (explained in Section 4.1), one may wonder whether this exclusion is coincidental with the dimension $d=1$. In this subsection, we discuss the extension of our result to general bounded domains in $\mathbb{R}^{d}$ (provided an a-priori estimate on the regularity of the eigenfunctions of the Dirichlet Laplacian), where the result (Theorem 4.1) still holds only for $\gamma$ in the complement of the $L^{2}$-regime. We strive for brevity here, as we believe that the main open problem of interest raised by this article is to address all $\gamma>0$.
Consider a bounded domain $D\subset\mathbb{R}^{d}$ with $d\geq 2$ (in $d=1$, $D$ must be an interval, and so the situation is easily seen to be equivalent to taking $D=[0,1]$). Let $\lambda_{k}$ be the $k^{\mathrm{th}}$ smallest eigenvalue (with multiplicity) of $-\Delta$ in $D$ with Dirichlet boundary conditions, and let $e_{k}$ denote a corresponding eigenfunction, normalized so that $\|e_{k}\|_{L^{2}(D)}=1$. Let $g:=(g_{k})_{k\in\mathbb{N}}$ denote a sequence of i.i.d. standard Gaussian random variables, and consider the Gaussian Fourier series
@@
This object is also known as the log-correlated Gaussian field in $D$ [LGF-survey, Proposition 1.1(3)] [^4] To match the description in [LGF-survey, Proposition 1.1(3)], note that, as explained in [FGF-survey, Section 9], the set $\{\phi_{k}:=\lambda_{k}^{-d/4}e_{k}\}_{k\in\mathbb{N}}$ forms an orthonormal basis of the Hilbert space that is the closure of Schwartz functions with respect to the inner-product $\langle f,g\rangle_{d}:=\langle(-\Delta)^{d/4}f,(-\Delta)^{d/4}g\rangle_{L^{2}(D)}$.. One can then consider the associated GMC measure $\mu_{\gamma,g}^{D}$, as well as the multiplicative chaos associated to the non-Gaussian Fourier series
@@
where the $a_{k}$ again are i.i.d. satisfying (1.2). The normalization constant and the range $\lambda_{k}\leq 2^{2n}$ are taken so that
$\operatorname{Cov}\big{(}\mathcal{S}_{n}(x),\mathcal{S}_{n}(y)\big{)}=\log\Big{(}\frac{1}{d(x,y)}\wedge 2^{n}\Big{)}+O(1)\,,$ (4.2)
where the $O(1)$ holds uniformly over all $x,y\in D$ [^5] While this formula seems to be forklore as we could not find it explicitly in the literature, we note that one can follow the proof of [cgff, Theorem 3], which addresses the $d=2$ case and is written in that paper’s Section 4, to obtain the result in any $d\geq 2$.. In this form, we can recognize that the critical parameter for the associated GMC is $\gamma=\sqrt{2d}$ and the $L^{2}$-regime is $(0,\sqrt{d})$.
Let $\mu_{\gamma,a}^{D}$ denote the multiplicative chaos associated to the field $(\mathcal{S}_{n,a}(x))_{x\in D,n\in\mathbb{N}}$. Note that the non-triviality of $\mu_{\gamma,a}^{D}$ for all $\gamma\in(0,\sqrt{2d})$ follows from the main result of [Junnila] only if the conditions of [Junnila, Definition 4] are satisfied by the field; in particular, [Junnila, Equation (5)] imposes certain regularity conditions on the $e_{k}$.
Similarly, so long as the $e_{k}$ satisfy certain regularity conditions, our proof gives mutual absolute continuity between $\mu_{\gamma,a}^{D}$ and a coupled GMC measure $\mu_{\gamma,g}^{D}$ outside of the $L^{2}$-regime.
@@
We now briefly describe how the proof of Theorem 4.1 differs from the proof of 1.1 with the aim of explaining the regularity condition (4.3).
The proof of 1.1 was outlined in Section 1.3, and we use the terminology of this outline here. Condition (4.3) is needed for exactly one step: the discretization step (we are able to move to a discrete hierarchical field by using the regularity of the eigenfunctions). We seek to define a nested sequence of partitions of $D$ as well as an approximation of $\mathcal{S}_{n,a}$ that is piecewise constant on each piece of the $n^{\mathrm{th}}$ partition, so that the approximation can be considered as a tree-indexed discrete field. Define the increments
@@
Thus, $\mathcal{X}_{n}(x)-\mathcal{X}_{n}(y)$ becomes small once $\|x-y\|\ll\tau_{n}^{-1/d}$. From this heuristic calculation, we see that, in order to make the chaining argument underlying Proposition 2.3 work, the $n^{\mathrm{th}}$ partition in our nested sequence of partitions of $D$ should contain (slightly more than) $\tau_{n}$ pieces of diameter (slightly less than) $\tau_{n}^{-1/d}$; in particular, in analogy with the bound on $T_{n}$ from (2.1), $n^{4d}\tau_{n}$ pieces in the $n^{\mathrm{th}}$ partition will suffice. Thus, following the proof of Proposition 2.3 will allow us to move to a discrete model on a tree where the set of particles $\mathcal{N}_{n}$ in the $n^{\mathrm{th}}$ generation satisfies $|\mathcal{N}_{n}|\leq n^{4d}\tau_{n}$.
Next is the coupling step. Following the proof of Lemma 2.5 with the above bound on $|\mathcal{N}_{n}|$, the Yurinskii coupling (Lemma 2.4) yields that the right-hand of (2.20) should be replaced by
@@
After the discretization and the coupling step, the proof proceeds exactly as in the proof of 1.1, since we have moved completely to the comparison of two chaoses associated to discrete models on the same tree (i.e., the space $D$ is no longer relevant).
@@ Acknowledgements
Both authors thank Ofer Zeitouni for suggesting the problem and for many helpful discussions. YHK thanks Klara Courteaut for a helpful discussion on the log-correlated Gaussian field in general domains and for pointing us to [cgff]. XK was supported by Israel Science Foundation grant number 615/24.
@@ Appendix A Chaos measures are supported on their thick points
In this section we provide a proof for Lemma 3.1. This is not novel, see for example [RV14, Theorem 4.1] for a similar argument in the Gaussian case. However, we did not find the result written down explicitly in the non-Gaussian case.
@@
$E_{n,j}^{+}$ :=$:=\left\{t\in[0,1]:\widetilde{S}_{n,\widetilde{a}}(t)>(\gamma+1/j)\log(2)n\right\}\,,$
$E_{n,j}^{-}$ :=$:=\left\{t\in[0,1]:\widetilde{S}_{n,\widetilde{a}}(t)<(\gamma-1/j)\log(2)n\right\}\,.$
@@
$\mathcal{T}^{c}\subseteq\bigcup_{j\in\mathbb{N}}\limsup_{n\to\infty}E_{n,j}^{+}\cup E_{n,j}^{-}\,.$
Thus, it is enough to show that $\mathbb{P}$-a.s. $\widetilde{\mu}_{\gamma,\widetilde{a}}[\limsup_{n\to\infty}E_{n,j}^{+}\cup E_{n,j}^{-}]=0$ for all $j\in\mathbb{N}$. We show this using the Borel-Cantelli lemma for the measure $Q$ on $\Omega\times[0,1]$ defined by $Q[A]:=\mathbb{E}[\int\mathbbm{1}_{A}\widetilde{\mu}_{\gamma,\widetilde{a}}(dt)]$, so that it suffices to show that for all $j\in\mathbb{N}$,
@@
together with Lemma 2.2 to obtain
@@
Combining (A.3), (A.4) yields (A.1). That $\widetilde{\mu}_{\gamma,\widetilde{g}}[\mathcal{T}^{c}]=0$ follows by taking $\widetilde{a}:=\widetilde{g}$ above. ∎
@@ Appendix B Properties of our partitions
We first prove (B.1). By construction, we have
[…]
The bound on $|I_{n}(w)|$ provided by Eq. 2.1 yields (B.1). We show (B.2) similarly: by construction,
[…]

### v2

@@ 4. Generalizations and Open Questions
Here, we discuss some natural extensions and open questions raised by our work. To explain what our proof strategy can and can not handle, we frequently refer to the proof outline in Section˜1.3. In particular, recall from there the “discretization step” and “coupling step.”
@@ 4.1. Critical chaos
The construction of the critical multiplicative chaos has not yet appeared for the random Fourier series (1.1); indeed, it is left as an open question (and for a more general class of non-Gaussian fields) in [25, Section 3.3]. Both the construction of the critical multiplicative chaos as well as absolute continuity with respect to a coupled critical GMC will be the subject of future work.
@@ 4.2. Extension to general bounded domains
In this article, we considered the space $D=[0,1]$ for convenience. In this subsection, we discuss the extension of our result to general bounded domains in $\mathbb{R}^{d}$ (provided an a-priori estimate on the regularity of the eigenfunctions of the Dirichlet Laplacian), where the result (Theorem˜4.1) still holds only for $\gamma$ in the complement of the $L^{2}$-regime.
Consider a bounded domain $D\subset\mathbb{R}^{d}$ with $d\geq 2$ (in $d=1$, $D$ must be an interval, and so the situation is easily seen to be equivalent to taking $D=[0,1]$). Let $\lambda_{k}$ be the $k^{\mathrm{th}}$ smallest eigenvalue (with multiplicity) of $\Delta$ in $D$ with Dirichlet boundary conditions, and let $e_{k}$ denote a corresponding eigenfunction, normalized so that $\|e_{k}\|_{L^{2}(D)}=1$. Let $g:=(g_{k})_{k\in\mathbb{N}}$ denote a sequence of i.i.d. standard Gaussian random variables, and consider
@@
This object is also known as the log-correlated Gaussian field in $D$ [17, Proposition 1.1(3)] [^3] To match the description in [17, Proposition 1.1(3)], note that, as explained in [32, Section 9], the set $\{\phi_{k}:=\lambda_{k}^{-d/4}e_{k}\}_{k\in\mathbb{N}}$ forms an orthonormal basis of the Hilbert space that is the closure of Schwartz functions with respect to the inner-product $\langle f,g\rangle_{d}:=\langle(-\Delta)^{d/4}f,(-\Delta)^{d/4}g\rangle_{L^{2}(D)}$.. One can then consider the associated GMC measure $\mu_{\gamma,g}^{D}$, as well as the multiplicative chaos associated to the non-Gaussian log-correlated random Fourier series/random wave model
@@
where the $a_{k}$ again are i.i.d. satisfying Eq.˜1.2. The normalization constant and the range $\lambda_{k}\leq 2^{2n}$ are taken so that
$\operatorname{Cov}\big(\mathcal{S}_{n,a}(x),\mathcal{S}_{n,a}(y)\big)=\log\Big(\frac{1}{d(x,y)}\wedge 2^{n}\Big)+O(1)\,$, (4.2)
where the $O(1)$ holds uniformly over all $x,y\in D$ [^4] While this formula seems to be forklore as we could not find it explicitly in the literature, we note that one can follow the proof of [39, Theorem 3], which addresses the $d=2$ case and is written in that paper’s Section 4, to obtain the result in any $d\geq 2$.. In this form, we can recognize that the critical parameter for the associated GMC is $\gamma=\sqrt{2d}$ and the $L^{2}$-regime is $(0,\sqrt{d})$.
Let $\mu_{\gamma,a}^{D}$ denote the multiplicative chaos associated to the field $(\mathcal{S}_{n,a}(x))_{x\in D,n\in\mathbb{N}}$. Note that the non-triviality of $\mu_{\gamma,a}^{D}$ for all $\gamma\in(0,\sqrt{2d})$ follows from the main result of [25] only if the conditions of [25, Definition 4] are satisfied by the field; in particular, [25, Eq. (5)] imposes certain regularity conditions on the $e_{k}$.
Similarly, as long as the $e_{k}$ satisfy certain regularity conditions, our proof gives mutual absolute continuity between $\mu_{\gamma,a}^{D}$ and a coupled GMC measure $\mu_{\gamma,g}^{D}$ outside of the $L^{2}$-regime.
@@
Remark 4.2.
Similarly to the outline in Section 1.3, we demonstrate below thats the only step which needs $\gamma>\sqrt{d}$ in the proof of Theorem 4.1 is the coupling step using the Yurinskii coupling (Lemma˜2.5). In [14] they explain how the high dimensional coupling they prove suffices to show that for all $d\geq 2$ one can couple $\mu_{\gamma,g}^{D}$ and $\mu_{\gamma,a}^{D}$ to be almost surely mutually absolutely continuous when the domain is taken to be $D=(0,1]^{d}$. The restriction on the domain comes from the fact that their CLT is not covariance agnostic, but requires control on the spectral norm of the covariance matrix of $(\widetilde{a}_{n}(v))_{v\in\mathcal{K}}$ from Eq.˜2.4, where $\mathcal{K}\subseteq\{1,\dots,T_{n}\}$ is arbitrary with $|\mathcal{K}|\leq 2^{n}$. Their estimate on the spectral norm crucially uses that the eigenfunctions of the Dirichlet Laplacian on unit cubes are given by products of signs and cosines. Thus, extending Theorem 4.1 to all subcritical $\gamma$ in the full generality of domains considered remains open.
We now briefly describe how the proof of Theorem˜4.1 differs from the proof of ˜1.1 with the aim of explaining the regularity condition (4.3).
The proof of ˜1.1 was outlined in Section˜1.3, and we use the terminology of this outline here. Condition (4.3) is needed for exactly one step: the discretization step (we are able to move to a discrete hierarchical field by using the regularity of the eigenfunctions). We seek to define a nested sequence of partitions of $D$ as well as an approximation of $\mathcal{S}_{n,a}$ that is piecewise constant on each piece of the $n^{\mathrm{th}}$ partition, so that the approximation can be considered as a tree-indexed discrete field. Define the increments
@@
Thus, $\mathcal{X}_{n}(x)-\mathcal{X}_{n}(y)$ becomes small once $\|x-y\|\ll\tau_{n}^{-1/d}$. From this heuristic calculation, we see that, in order to make the chaining argument underlying Proposition˜2.3 work, the $n^{\mathrm{th}}$ partition in our nested sequence of partitions of $D$ should contain (slightly more than) $\tau_{n}$ pieces of diameter (slightly less than) $\tau_{n}^{-1/d}$; in particular, in analogy with the bound on $T_{n}$ from Eq.˜2.1, $n^{4d}\tau_{n}$ pieces in the $n^{\mathrm{th}}$ partition will suffice. Thus, following the proof of Proposition˜2.3 allows us to move to a discrete model on a tree where the set of particles $\mathcal{N}_{n}$ in the $n^{\mathrm{th}}$ generation satisfies $|\mathcal{N}_{n}|\leq n^{4d}\tau_{n}$.
Next is the coupling step. Following the proof of Lemma˜2.5 with the above bound on $|\mathcal{N}_{n}|$, the Yurinskii coupling (Lemma˜2.4) yields that the right-hand of Eq.˜2.20 should be replaced by
@@
After the discretization and the coupling step, the proof proceeds exactly as in the proof of ˜1.1, since we have moved completely to the comparison of two chaoses associated to discrete models on the same tree (i.e., the space $D$ is no longer relevant).
@@ Acknowledgements
Both authors thank Ofer Zeitouni for suggesting the problem and for many helpful discussions. YHK thanks Klara Courteaut for a helpful discussion on the log-correlated Gaussian field in general domains and for pointing us to [39]. We thank the anonymous referee for a careful reading of the paper and for comments and detailed suggestions which helped us to improve the manuscript.
@@ Appendix A Chaos measures are supported on their thick points
In this section we provide a proof for Lemma 3.1. This is not novel, see for example [38, Theorem 4.1] for a similar argument in the Gaussian case. However, we did not find the result written down explicitly in the non-Gaussian case.
@@
$E_{n,j}^{+}$ $:=\left\{t\in[0,1]:\widetilde{S}_{n,\widetilde{a}}(t)>(\gamma+1/j)\log(2)n\right\}\,$,
$E_{n,j}^{-}$ $:=\left\{t\in[0,1]:\widetilde{S}_{n,\widetilde{a}}(t)<(\gamma-1/j)\log(2)n\right\}\,.$
@@
$\mathcal{T}^{c}\subseteq\bigcup_{j\in\mathbb{N}}\limsup_{n\to\infty}(E_{n,j}^{+}\cup E_{n,j}^{-})\,.$
Thus, it is enough to show that $\mathbb{P}$-a.s. $\widetilde{\mu}_{\gamma,\widetilde{a}}[\limsup_{n\to\infty}(E_{n,j}^{+}\cup E_{n,j}^{-})]=0$ for all $j\in\mathbb{N}$. We show this using the Borel-Cantelli lemma for the measure $Q$ on $\Omega\times[0,1]$ defined by $Q[A]:=\mathbb{E}[\int\mathbbm{1}_{A}\widetilde{\mu}_{\gamma,\widetilde{a}}(dt)]$, so that it suffices to show that for all $j\in\mathbb{N}$,
@@
together with Lemma˜2.2 to obtain
@@
Combining Eq.˜A.3, Eq.˜A.4 yields Eq.˜A.1. That $\widetilde{\mu}_{\gamma,\widetilde{g}}[\mathcal{T}^{c}]=0$ follows by taking $\widetilde{a}:=\widetilde{g}$ above. ∎
@@ Appendix B Properties of our partitions
We first prove Eq.˜B.1. By construction, we have
$|\mathcal{N}_{n+h}(w)|\leq\sum_{k=n+1}^{n+h}|I_{n}(w)|\,2^{k}f(k)\leq|I_{n}(w)|2^{n+h+1}f(n+h)\,.$
The bound on $|I_{n}(w)|$ provided by Eq.˜2.1 yields Eq.˜B.1. We show Eq.˜B.2 similarly:
$|\mathcal{D}_{I,n}(v)|\leq\sum_{k=1}^{n}|I+t_{n}(v)|\,2^{k}f(k)\leq|I|\,2^{n+1}f(n)=T_{n}\,|I|\,.\qed$